# Towards Robust Stability Prediction in Smart Grids: GAN-based Approach under Data Constraints and Adversarial Challenges

### v1

@@ VI-D Benchmarking Against State-of-the-Art
ANN [1] 0.996
XGBoost [3] 0.994
CNN [37] 0.870
LSTM [14] 0.990
ANN [35] 0.973
Decision Tree [3] 0.974
KNN [3] 0.875
GAN-Stability 0.975 0.989111Whitebox - 0.955222Greybox
11footnotetext: Adversarial robustness measured under whitebox conditions.22footnotetext: Adversarial robustness measured under greybox conditions.
VI-E Training Time
@@
VI-F Hardware Implementation
In order to demonstrate the practicality of our framework in real-world applications, we conducted an experiment using a Raspberry Pi 4 Model B, which features a 4GB RAM and a Quad-core Cortex-A72 (ARM v8) 64-bit SoC running at 1.5GHz, Raspberry Pi OS 12 (Debian Bookworm porting) [55]. We used Python 3.11.2 and Pytorch 2.4.1. This configuration represents one of the most basic and cost-effective hardware setups available on the market. The complete results of our experiment can be found in Table V.
TABLE V: Time results (in seconds) for training of one epoch and testing for one batch for our models in a Raspberry Pi.
@@ VI-F Hardware Implementation
VII conclusion
Our paper presents a pioneering framework using a GAN model to predict smart grid stability, effectively addressing the challenge of limited data accessibility by focusing on stable instances from available datasets. This approach aligns with real-world constraints where instability data is scarce. Our model achieves a commendable accuracy of 0.975 in stability prediction, even without instability instances in the training data, using just 32.85% of the dataset. Additionally, by incorporating a novel adversarial training layer, our framework demonstrates robustness against state-of-the-art adversarial attacks in both whitebox and greybox scenarios, classifying these attacks as instances of instability. We benchmarked our model against state-of-the-art approaches and found that while it may not surpass current best performances, it still offers high accuracy and enhanced robustness, a feature not commonly found in existing literature. In summary, our study advances stability prediction for smart grids by offering high accuracy and resilience, addressing data scarcity challenges, and enhancing operational efficiency and reliability in evolving energy landscapes with emerging cybersecurity threats.

### v2

@@ 6.4 Benchmarking Against State-of-the-Art
ANN [1] 0.996 N/A
ANN [62] 0.985 N/A
XGBoost [3] 0.994 N/A
CatBoost [37] 0.996 N/A
CNN [38] 0.870 N/A
LSTM [14] 0.990 N/A
ANN [36] 0.973 N/A
Decision Tree [3] 0.974 N/A
KNN [3] 0.875 N/A
GAN-Stability 0.981 0.9891 - 0.9882
  • 1
    Adversarial robustness measured under white-box conditions.
  • 2
    Adversarial robustness measured under grey-box conditions.
6.5 Training Time
@@
6.6 Hardware Implementation
In order to demonstrate the practicality of our framework in real-world applications, we conducted an experiment using a Raspberry Pi 4 Model B, which features a 4GB RAM and a Quad-core Cortex-A72 (ARM v8) 64-bit SoC running at 1.5GHz, Raspberry Pi OS 12 (Debian Bookworm porting) [63]. We used Python 3.11.2 and Pytorch 2.4.1. This configuration represents one of the most basic and cost-effective hardware setups available on the market. The complete results of our experiment can be found in Table 5.
Table 5: Time results (in seconds) for training of one epoch and testing for one batch for our models in a Raspberry Pi.
@@ 6.6 Hardware Implementation
7 Limitation
In this section, we outline some of the limitations of this study.
  • 1.
    Real-World Deployment Constraints: While the model demonstrated efficient real-time decision-making on a single-board computer, real-world smart grid environments may introduce additional challenges, such as communication latency, hardware limitations, and integration complexities. Deploying the proposed model requires seamless compatibility with existing grid monitoring, control, and communication systems, including Supervisory Control and Data Acquisition (SCADA) systems, phasor measurement units (PMUs), and energy management systems (EMS). Achieving this integration may necessitate modifications or middleware solutions.
  • 2.
    Lack of Explicit Interpretability: GAN-based models are inherently complex and may lack transparency, making it difficult for operators to understand the reasoning behind specific instability detections.
  • 3.
    Requirement for a Comprehensive Stable-Class Dataset: The effectiveness of this approach depends on access to a diverse and comprehensive dataset of stable-class instances. If the dataset lacks sufficient variability, the model may struggle to distinguish true instability from rare but valid variations within stable conditions.
  • 4.
    Evaluation on Larger Datasets: While the model has been tested on a dataset containing 60000 stable and unstable samples, its performance on significantly larger datasets with greater variability due to the lack of publicly available large-scale smart grid dataset remains untested. Further validation on diverse grid configurations, renewable energy penetration levels, is necessary to fully assess its scalability and robustness.
8 Conclusion
Our paper presents a pioneering framework using a GAN model to predict smart grid stability, effectively addressing the challenge of limited data accessibility by focusing on stable instances from available datasets. This approach aligns with real-world constraints where instability data is scarce. Our model achieves a commendable accuracy of 0.981 in stability prediction, even without instability instances in the training data, using just 32.85% of the dataset. Additionally, by incorporating a novel adversarial training layer, our framework demonstrates robustness against state-of-the-art adversarial attacks in both whitebox and greybox scenarios, classifying these attacks as instances of instability. We benchmarked our model against state-of-the-art approaches and found that while it may not surpass current best performances, it still offers high accuracy and enhanced robustness, a feature not commonly found in existing literature. In summary, our study advances stability prediction for smart grids by offering high accuracy and resilience, addressing data scarcity challenges, and enhancing operational efficiency and reliability in evolving energy landscapes with emerging cybersecurity threats.
8.1 Future Work
While our framework demonstrates strong performance in stability prediction and adversarial robustness, several directions remain for future exploration. First, we plan to expand the model’s capabilities by incorporating real-world instability instances as they become available. This would enable direct training on both stable and unstable samples, potentially through continual learning techniques. Second, we seek to enhance the diversity and semantic quality of the generated OOD samples. Approaches may include conditioning the generator on domain-specific attributes or employing contrastive objectives to promote clearer separation between stable and unstable regions in the latent space. Another promising direction involves integrating domain knowledge and physics-informed constraints into the model to ensure interpretability and alignment with the underlying dynamics of power grids. Finally, we are exploring deployment in real-time grid monitoring environments. This will require adapting the model to meet strict latency constraints, process streaming data, and maintain robustness over extended operational periods without retraining. Additionally, we are considering evaluation with historical SCADA/PMU data to assess the feasibility of a real-world application.